OpenStudy (anonymous):

Devon and his friends bought strawberry wafers for $3 per packet and chocolate wafers for $1 per packet at a carnival. They spent a total of $30 to buy a total of 22 packets of wafers of the two varieties. Part A: Write a system of equations that can be solved to find the number of packets of strawberry wafers and the number of packets of chocolate wafers that Devon and his friends bought at the carnival. Define the variables used in the equations.

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

do you know what a system of equations is?

OpenStudy (anonymous):

not really... like, a couple of equations?

OpenStudy (anonymous):

well sort of

OpenStudy (anonymous):

the goal of this question is to write one....... so there is 2 equations in a system of equation, they are used to find cost of something like with 2 variables.

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so first you need to determine 2 variables(the numbers you are tying to find in the problem)

OpenStudy (anonymous):

do know which two?

OpenStudy (anonymous):

s and c

OpenStudy (anonymous):

strawberry for s chocolate for c right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now we need to make 2 equations to find how many of each wafers Devon bought

OpenStudy (anonymous):

thats the part that idk

OpenStudy (anonymous):

s + c = 22 he bought a total of 22 wafers. 3s + c = 30 amount of money spent and how much each wafer is

OpenStudy (anonymous):

does this make sense? that will be it for this question since it doesn't ask you to solve...

OpenStudy (anonymous):

oooo yes!! and, part b asks me to solve..

OpenStudy (anonymous):

Part B: How many packets of chocolate wafers and strawberry wafers did they buy? Explain how you got the answer and why you selected a particular method to get the answer.

OpenStudy (anonymous):

well at least part a is done :D

OpenStudy (anonymous):

can you help with part b?

OpenStudy (anonymous):

ok sure what don't you get?

OpenStudy (anonymous):

how many of each

OpenStudy (anonymous):

ok so first lets set up the equation, put it on of top of another like this s + c = 22 3s + c = 30

OpenStudy (anonymous):

now we have to isolate one variable

OpenStudy (anonymous):

in this case, c is alone for both equations so we can subtract it

OpenStudy (anonymous):

correct..

OpenStudy (anonymous):

3s + c = 30 - s + c = 22 ---------------- 2s = 8

OpenStudy (anonymous):

still following?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so there is 4 strawberry wafers, now we just plug in or substitute in 4 for s in which ever equation you want lets do s + c = 22 since its easier 4 + c = 22 c = 18

OpenStudy (anonymous):

now check your answer in the other equation, 3s + c = 30 3(4) + 18 = 30 12 + 18 = 30 correct!!

OpenStudy (anonymous):

oo!! Thanks so much!!

OpenStudy (anonymous):

no problem